what is the property term for 8/8=1? 5/?=5
Answer by Mathtut(3670)

Multiplicative Identity Property: The product of any number and one is that number. For example 5 * 1 = 5
:
and any number divided by itself is 1
:
in this case ?=1.....5/5=1
:
remember multiplication and division are inverses of each other
:
so 5*1=5 and 5/5=1
:
8/8=1 or 8/1=8